Find the periodic withdrawals PMT for the annuity given. Round your answer to the nearest cent
$450,000 at 2%, paid out monthly for 18 years
PMT = $ ?

Answer by Theo(13342)

if you withdraw the money at the end of the month, then you need this formula.

ANNUITY FOR A PRESENT AMOUNT WITH END OF TIME PERIOD PAYMENTS

if you withdraw the money at the beginning of the month, then you need this formula.

ANNUITY FOR PRESENT AMOUNT WITH BEGINNING OF TIME PERIOD PAYMENTS

both these formulas are found in the following link.

https://www.algebra.com/algebra/homework/Finance/THEO-2016-04-29.lesson#formulas

you can do it manually or you can use the following online calculator that will do it for you.

http://arachnoid.com/finance/index.html

i used the online calculator.

i did end of time period first and then beginning of time period.
i also did 2% per year and 2% per month because it wasn't clear whether you meant at 2% per year compounded monthly or at 2% per month.

following are the results.

first is end of month withdrawals at 2% per year compounded monthly.
withdrawal = 2482.50 per month.
$$$

second is beginning of month withdrawals at 2% per year compounded monthly.
withdrawal = 2478.37 per month.
$$$

third is end of month withdrawals at 2% per month.
withdrawal = 9126.67 per month.
$$$

fourth is beginning of month withdrawals at 2% per month.
withdrawal = 8947.72
$$$

you enter present value = 450,000
you enter future value = 0
you enter number of time periods = 18 * 12 = 216
you enter interest rate percent = 2/12 = .1666667 if 2% per year compounded monthly.
you enter interest rate percent = 2 if 2% per monthy.
you select payment at beginning or end of time period.
you click on pmt
you get your answer.